# Emergent Quantum Walk Dynamics from Classical Interacting Particles

### v3

@@ Abstract
$R_{x,s}=\frac{1}{2}(1+\frac{s\delta_{x}}{\lvert\delta_{x}\rvert})\frac{\delta_{x}}{n_{x}^{s}},$
Here $\delta_{x}=\tilde{n}_{x}^{+1}-n_{x}^{+1}$, where $\tilde{n}_{x}^{+1}$ is defined by the same expression as $\tilde{N}_{0}$ in Eq. (35), with the local variables $n_{x}^{+1}$, $n_{x}^{-1}$, $\eta_{x}^{+1}$, and $\eta_{x}^{-1}$ playing the roles of $N_{0}$, $N_{1}$, $\phi_{0}$, and $\phi_{1}$, respectively.
@@
In this letter, we have developed classical interacting particle models to simulate the dynamics of 1D discrete-time quantum walks with a generalized coin. The classical models can be systematically generalized to higher-dimensional graphs. In the two-dimensional case, for example, the walker requires a $4\times 4$ coin operation acting on a two-qubit state, representing four possible directions: up, down, left, and right. Consequently, the box-ball model must be extended to include a transformation operation $\mathbb{T}$ that acts between four distinct boxes, with each site in the two-dimensional lattice hosting four boxes. By carefully extending these ideas, the box model can be generalized to arbitrary higher-dimensional graphs or even complex networks, where each node may be associated with multiple boxes corresponding to the available directions or connections. So, this approach offers a versatile and intuitive framework for simulating a broad range of quantum dynamical systems. Our future work will focus on formalizing this methodology and exploring its application to increasingly complex topologies.
@@
Quantum walks have found successful applications across a wide range of domains [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], from quantum search algorithms [34] to simulating relativistic quantum dynamics [41, 42] and modeling quantum active particles [46]. The present study establishes a connection between the generalised active spin model and all these domains where quantum walks are effectively applied.
@@
In the literature, numerous efforts have been undertaken to simulate quantum behavior using classical frameworks. Broadly, these attempts can be categorized into two approaches: in one, wave functions guide particle trajectories [9]; in the other, wave functions play no dynamical role [10, 11, 12, 13, 14, 15, 49]. Our approach aligns with the latter. The simplicity and accessibility of the framework presented in this study may offer a promising contribution toward foundational insights into the deeper structures underlying quantum theory.

### v4

# arXiv:2506.06896v4[quant-ph] 02 Mar 2026

@@ Abstract
In the box–ball formulation, essentially the boxes interact through the exchange of balls. This idea of ball relocation can be extended to the level of individual particle dynamics by employing an active spin model.
@@
$R_{x,s}=\frac{s}{2}(1+\frac{s\delta_{x}}{\lvert\delta_{x}\rvert})\frac{\delta_{x}}{n_{x}^{s}},$
Here $\delta_{x}=n_{x}^{+1}-\tilde{n}_{x}^{+1}$, where $\tilde{n}_{x}^{+1}$ is defined by the same expression as $\tilde{N}_{0}$ in Eq. (35), with the local variables $n_{x}^{+1}$, $n_{x}^{-1}$, $\eta_{x}^{+1}$, and $\eta_{x}^{-1}$ playing the roles of $N_{0}$, $N_{1}$, $\phi_{0}$, and $\phi_{1}$, respectively.
@@
In this letter, we develop classical interacting particle models that simulate one-dimensional discrete-time quantum walks with a generalized coin. The framework can be naturally extended to higher-dimensional lattices and complex networks by associating multiple boxes (directions) with each node and introducing appropriate transformation and shift operations.
@@
Quantum walks have found successful applications across a wide range of domains [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], from quantum search algorithms [34] to simulating relativistic quantum dynamics [41, 42] and modeling quantum active particles [46].
The present study establishes a connection between the generalised active spin model and all these domains where quantum walks are effectively applied.
@@
Conceptually, our approach belongs to the class of classical frameworks that reproduce quantum behavior without assigning a dynamical role to the wave function [10, 11, 12, 13, 14, 15, 49], providing a simple and accessible route toward deeper structural insights into quantum theory.